OpenStudy (anonymous):

in the equation ax+by=20 and bx+ay=16, the value for x=2 and y=1. What are the coefficient a and b?

OpenStudy (anonymous):

if you substitute the values then the equations become 2a+b=20 and 2b+a=16

OpenStudy (anonymous):

So try solving it or I can help you solve it

OpenStudy (anonymous):

yeah I know you get those but what next

OpenStudy (anonymous):

okay so multiply the first equation by 2 you get 4a+ 2b = 40

OpenStudy (anonymous):

the 2nd equation is 2b+a = 16 so now subtract this equation from 4a+2b = 40

OpenStudy (anonymous):

you will get (4a+2b)-(a+2b) = 40-16

OpenStudy (anonymous):

From this you will get the value of a

OpenStudy (anonymous):

Using the value of a you will get the value of b

OpenStudy (anonymous):

well 40- 16= 24 so then you know that a=24 would you substitute 24 for a in the next equation?

OpenStudy (anonymous):

Nope see on the left hand side of the equation, it is 3a not a

OpenStudy (anonymous):

So 3a = 24

OpenStudy (anonymous):

ok...so then how would it look? don't solve the equation but just so I can see how it will be set up

OpenStudy (anonymous):

i will set it up in a drawing and attach it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

so then it would look like 2(24)+b=20